OpenStudy (anonymous):

Multiplying Polynomial 20ab^a.16a^13b^15 I have solved it and this was my final answer Just want to be sure if I got it : 320ab^a a^13b^15

OpenStudy (ranga):

\[\Large (20ab ^{a})\times (16a ^{13}b ^{15})\] Is that how the question looks?

OpenStudy (anonymous):

No Parenthesis

OpenStudy (anonymous):

\[20ab ^{a}.16a ^{13}b ^{15}\]

OpenStudy (ranga):

20 x 16 = 320 a x a^13 = ? b^a x b^15 = ?

OpenStudy (ranga):

\[\Large a ^{m} \times a ^{n} = a ^{m+n}\]

OpenStudy (ranga):

Do you follow?

OpenStudy (anonymous):

Yes so My answer is correct then

OpenStudy (ranga):

No. You need to simplify further.

OpenStudy (anonymous):

If I simplify further 320a\[320a ^{14}b ^{a+15}\]

OpenStudy (ranga):

Good!

OpenStudy (anonymous):

Thank you so much

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

I have another question but I didnt know how to do that you might have to solved it for me then next time I can do it myself by following your steps